### v2

@@ 6.2 Stochastic Lotka-Volterra
We first compare the accuracy of the MAP points, posterior samples, and posterior predictives across the three approaches. The MAP points are obtained using the same method as in Section 6.2. For posterior sampling, we first sample a “ground truth” ${\bf x}^{*}\sim\pi({\bf x})$ and obtain the associated ground truth reduced observation ${\bf y}^{*}_{\rm proj}={\bf V}_{\rm proj}^{\top}(\Psi({\bf x}^{*})+0.25{\boldsymbol{\epsilon}})$. Then, we use the three approaches to sample from the posterior $\pi_{\rm proj}({\bf x}_{\rm proj}|{\bf y}^{*}_{\rm proj})$. This allows us to obtain approximate samples ${\bf x}\sim\pi({\bf x}|{\bf y}^{*})$. The posterior predictives are obtained by solving the forward model for the generated parameters. Through Fig. 6, we observe that the MAP points, posterior samples, and predictives produced by PCP-Map and COT-Flow are more concentrated around the ground truth than those produced by NPE.
Figure 6: Prior, MAP point, posterior samples, and predictives quality comparisons between PCP-Map, COT-Flow, and NPE. Each Row: left: prior (first row) or posterior samples in gray and ${\bf x}^{*}$ in black. middle: 2D image of the wave amplitudes simulated using ${\bf x}^{*}$ (first row) or posterior samples for 100 time grids. right: wave amplitudes simulated using 50 prior (first row) or posterior samples at t = 22, 69, 94. Here ${\bf y}^{*}$ is plotted in black.
@@
We perform the simulation-based calibration (SBC) analysis described in  [48, App. D.2] to further assess the three approaches’ accuracy; see Fig. 7. We can see that, while they are all well calibrated, the cumulative density functions of the rank statistics produced by PCP-Map align almost perfectly with the CDF of a uniform distribution besides a few outliers.
@@
Figure 7: SBC Analysis for PCP-Map, COT-Flow and NPE. Each colored line represents the empirical cumulative density function (CDF) of the SBC rank associated with one posterior sample dimension.
@@
Finally, we analyze the three approaches’ efficiency in terms of the number of forward model evaluations. We train the models using the best hyperparameter combinations from the pilot run on two extra datasets with 50k and 20k samples. We compare the posterior samples’ mean and standard deviation against ${\bf x}^{*}$ across three approaches trained on the 100k, 50k, and 20k sized datasets as presented in Fig. 8. We see that PCP-Map and COT-Flow can generate posterior samples centered more closely around the ground truth than NPE using only 50k training samples, which translates to higher computational efficiency since fewer forward model evaluations are required.
Figure 8: Posterior sample quality comparisons between PCP-Map, COT-Flow, and NPE trained using 20k, 50k, and 100k samples. We report the relative normed error between the posterior sample mean (over 100 samples) and the ground truth parameter. The gray bands represent regions within one standard deviation of the means.
@@ 6.4 Comparing PCP-Map to amortized CP Flow
Table 7: Computational cost comparisons between amortized CP-Flow and PCP-Map in terms of GPU time in seconds (s). We report the mean and standard deviation of the training and validation mean NLL, the mean and standard deviation of the GPU times over five training runs, and the number of parameters in million(M), respectively.
@@ 7 Discussion
In our numerical experiments, the comparison to the SMC-ABC approach for the stochastic Lotka-Volterra problem illustrates common trade-offs when selecting conditional sampling approaches. Advantages of the ABC approach include its strong theoretical guarantees and well-known guidelines for choosing the involved hyper-parameters (annealing, burn-in, number of samples to skip to reduce correlation, etc.). The disadvantages are that ABC typically requires a large number of likelihood evaluations to produce (approximately) i.i.d. samples and low-variance estimators in high-dimensional parameter spaces; the computation is difficult to parallelize in the sequential Monte Carlo setting, and the sampling process is not amortized over the conditioning variable ${\bf y}^{*}$, i.e., it needs to be recomputed whenever ${\bf y}^{*}$ changes.
Comparisons to the flow-based NPE method [48] for the high-dimensional 1D shallow water equations problem Section 6.3 illustrate the superior accuracy achieved by our approaches. The results indicate that both methods can learn high-dimensional COT maps. Here, the number of effective parameters in the dataset was $n=14$, and the number of effective measurements was $m=3500$. Particularly worth noting is that PCP-Map, on average, converges in around 715 seconds on one GPU on this challenging high-dimensional problem. Moreover, when compared to the amortized CP-Flow approach, PCP-Map achieves significantly faster convergence while providing a working computational scheme to the static COT problem.
@@
Given the empirical nature of our study, we paid particular attention to the setup and reproducibility of our numerical experiments. To show the robustness of our approaches to hyperparameters and to provide guidelines for hyperparameter selection in future experiments, we report the results of a simple two-step heuristic that randomly samples hyperparameters and identifies the most promising configurations after a small number of training steps. We stress that the same search space of hyperparameters is used across all numerical experiments.
@@
Beyond the approaches surveyed and discussed, diffusion models [52] gained much success recently across many applications. In particular, following the procedure in [52, Appendix I.4], the resulting unconditional generator can be used to obtain various conditional distributions as long as there is a tractable and differentiable log-likelihood function; see, e.g., applications to image generation and time series imputation in [4, 54]. We omit discussions on diffusion models since the notion of measure transport is not immediately obvious in this setting as the reference samples are evolved stochastically. Besides diffusion models, flow matching has been recently proposed as an alternative training scheme for continuous normalizing flows [33] with connections to OT. Since flow matching alleviates the need for time integration, investigating extensions to COT merits further investigation [27].

### v3

@@ 6.2 Stochastic Lotka–Volterra
We first compare the accuracy of the MAP points, posterior samples, and posterior predictives across the three approaches. The MAP points are obtained using the same method as in Section 6.2. For posterior sampling, we first sample a “ground truth” ${\bf x}^{*}\sim\pi({\bf x})$ and obtain the associated ground truth reduced observation ${\bf y}^{*}_{\rm proj}={\bf V}_{\rm proj}^{\top}(\Psi({\bf x}^{*})+0.25{\boldsymbol{\epsilon}})$. Then, we use the three approaches to sample from the posterior $\pi_{\rm proj}({\bf x}_{\rm proj}|{\bf y}^{*}_{\rm proj})$. This allows us to obtain approximate samples ${\bf x}\sim\pi({\bf x}|{\bf y}^{*})$. The posterior predictives are obtained by solving the forward model for the generated parameters. Through Figure 5, we observe that the MAP points, posterior samples, and predictives produced by PCP-Map and COT-Flow are more concentrated around the ground truth than those produced by NPE.
Figure 5: Shallow water experiment, Section 6.3. Comparison of MAP estimation accuracy, posterior sampling quality, and posterior predictive accuracy among PCP-Map, COT-Flow, and NPE. Column 1: Ground truth parameter ${\bf x}^{\ast}$ (black), prior parameter samples (gray, row 1), generated posterior samples (gray, row 2-4), and MAP points (purple cross, row 2-4). Column 2: 2D wave simulation images over 100 time and space grids. Horizontal color lines mark 3 time slices shown in Columns 3–5. Ground truth simulation (row 1) and simulation from generated posterior samples (row 2-4). Columns 3–5: Wave amplitudes across the 3 time slices. Ground truth simulations (black), simulations from prior samples (row 1) and generated posterior samples (rows 2–4).
@@
We perform the simulation-based calibration (SBC) analysis described in  [50, App. D.2] to further assess the three approaches’ accuracy; see Figure 6. We can see that, while they are all well calibrated, the cumulative distribution functions of the rank statistics produced by PCP-Map align almost perfectly with the CDF of a uniform distribution besides a few outliers.
@@
Figure 6: Shallow water experiment, Section 6.3. SBC analysis for PCP-Map, COT-Flow, and NPE. Each colored line represents the empirical cumulative distribution function (CDF) of the SBC rank associated with one posterior sample dimension. The more the empirical CDF resembles the CDF of the uniform distribution (i.e., the identity function), the better.
@@
Finally, we analyze the three approaches’ efficiency in terms of the number of forward model evaluations. We train the models using the best hyperparameter combinations from the pilot run on two extra datasets with 50k and 20k samples. We compare the posterior samples’ mean and standard deviation against ${\bf x}^{*}$ across three approaches trained on the 100k, 50k, and 20k sized datasets as presented in Figure 7. We see that PCP-Map and COT-Flow can generate posterior samples centered more closely around the ground truth than NPE using only 50k training samples, which translates to higher computational efficiency since fewer forward model evaluations are required.
Figure 7: Shallow water experiment, Section 6.3. Posterior sample quality comparison between PCP-Map, COT-Flow, and NPE trained on 20k, 50k, and 100k samples. Each plot shows the ground truth ${\bf x}^{\ast}$ (black), posterior mean (color), and standard deviation (gray band) over 100 samples.
@@ 6.4 Comparing PCP-Map to Amortized CP Flow
Table 7: Computational cost comparison (Section 6.4) between amortized CP-Flow and PCP-Map. We report the mean and standard deviation of GPU time (s), training/validation NLL, and number of model parameters $M$ over five runs. The better result is highlighted in bold.
@@ 7 Discussion
In our numerical experiments, the comparison to the SMC-ABC approach for the stochastic Lotka–Volterra problem illustrates common trade-offs when selecting conditional sampling approaches. Advantages of the ABC approach include its strong theoretical guarantees and well-known guidelines for choosing the involved hyper-parameters (annealing, burn-in, number of samples to skip to reduce correlation, etc.). The disadvantages are that ABC typically requires a large number of likelihood evaluations to produce (approximately) i.i.d. samples and low-variance estimators in high-dimensional parameter spaces; the computation is difficult to parallelize in the sequential Monte Carlo setting, and the sampling process is not amortized over the conditioning variable ${\bf y}^{*}$, i.e., it needs to be recomputed whenever ${\bf y}^{*}$ changes.
Comparisons to the flow-based NPE method [50] for the high-dimensional 1D shallow water equations problem (Section 6.3) illustrate the superior accuracy achieved by our approaches. The results indicate that both methods can learn high-dimensional COT maps. Here, the number of effective parameters in the dataset was $n=14$, and the number of effective measurements was $m=3500$. Particularly worth noting is that PCP-Map, on average, converges in around 715 seconds on one GPU on this challenging high-dimensional problem. Moreover, when compared to the amortized CP-Flow approach, PCP-Map achieves significantly faster convergence while providing a working computational scheme to the static COT problem.
@@
Given the empirical nature of our study, we paid particular attention to the setup and reproducibility of our numerical experiments. To show the robustness of our approaches to hyperparameters and to provide guidelines for hyperparameter selection in future experiments, we report the results of a simple two-step heuristic that randomly samples hyperparameters and identifies the most promising configurations after a small number of training steps.
@@
Beyond the approaches surveyed and discussed, diffusion models [55] gained much success recently across many applications. In particular, following the procedure in [55, Appendix I.4], the resulting unconditional generator can be used to obtain various conditional distributions as long as there is a tractable and differentiable log-likelihood function; see, e.g., applications to image generation and time series imputation in [4, 57]. We omit discussions on diffusion models since the notion of measure transport is not immediately obvious in this setting as the reference samples are evolved stochastically. Besides diffusion models, flow matching has been recently proposed as an alternative training scheme for continuous normalizing flows [34] with connections to OT. Since flow matching alleviates the need for time integration, investigating extensions to COT merits further investigation [29].